# Dynamic asymptotic dimension growth for group actions and groupoids

### v1

@@ 3.2. The BLR condtions
Elements of $\textnormal{Prob}(V)$ can be written as $\mu=\sum\limits_{v\in V}t_{v}v$, where $\mu(v)=t_{v}$ in $[0,1]$ and $\sum\limits_{v\in V}t_{v}=1$. Write $\textnormal{Prob}_{n}(V)$ for the subset of $\textnormal{Prob}(V)$ consisting of measures supported on at most $n+1$ points. A simplicial complex is a subspace $C$ of some $\bigcup\limits_{n=0}^{\infty}\textnormal{Prob}_{n}(V)$ such that $V\subset C$ and
$\left\{\sum\limits_{v\in V}t_{v}v\in\bigcup\limits_{n=0}^{\infty}\textnormal{Prob}_{n}(V)\mid t_{v}\neq 0\text{ for }v\in\textnormal{supp }(\mu)\right\}\in C.$
@@
To describe the following lemma, we review the notion of a simplicial action. Let $\Gamma$ be a discrete group, and let $C$ be a simplicial complex with the vertex set $V$. A simplicial action $g$ of $\Gamma$ on $C$ is an action that is induced from some action of $\Gamma$ on $V$ via the formula
Clearly, a simplicial action is isometric. The next technical lemma will be used in the proof of Proposition 3.12 and Theorems 3.14 and 4.17.
@@
The $d$-BLR condition was first defined in Bartels’ paper [Bar16], then discussed in the paper [GWY17] to connect with the finite asymptotic dimension. Below, we will generalize it to the growth case.
@@
If $h\preceq x^{\alpha}(0<\alpha<\frac{1}{2})$, then we get an equivariant version of the $h$-BLR condition, which is associated with the asymptotic dimension growth of groups.
@@
    For any $R>0$, there exists a nondecreasing function $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ such that for the finite subset $E=\{\gamma\in\Gamma\mid\ell_{\Gamma}(\gamma)<R\}$ of $\Gamma$, there exists a $\Gamma$ equivaraint open cover $\mathcal{U}$ of $X\times\Gamma$ such that the following hold.
@@
This proposition can be proved in a similar way as shown in [Proposition 4.5, [GWY17]]. In the proof of $(1)$ implying $(2)$, there is no extra restrictions on the growth rate of $h$, while in the case of $(2)$ implying $(1)$, the function $h$ must grow slower than $\sqrt{x}$.
The following theorem shows that an $h$-BLR action has dynamic asymptotic dimension function at most $h$. Under the assumption of subexponential growth of $h$, Theorem 4.17 will give a converse in coarse geometry in the next section.
@@
Let $\Gamma\curvearrowright X$ be an action with $X$ compact, and let $\mathcal{F}$ be the family of finite subgroups of $\Gamma$. Let $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a nondecreasing function. If $\Gamma\curvearrowright X$ is an $h$-BLR action for $\mathcal{F}$, then it has dynamic asymptotic dimension function at most $h$.
@@
First, we construct a desired open cover of $X$. Assume that $\Gamma\curvearrowright X$ is an $h$-$BLR$ action for $\mathcal{F}$. Given $R>0$, there exists a nondecreasing function $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ such that for the finite subset $E=\{\gamma\in\Gamma\mid\ell(\gamma)<R\}$ of $\Gamma$ and $\varepsilon=\frac{1}{3}10^{-h(R)}$, there exist a $(\Gamma,\mathcal{F})$-complex $C$ of dimension at most $h(R)$ and a continuous $(E,\varepsilon)$-equivariant map $f:X\rightarrow C$. By Lemma 3.9, we may assume that there exists a finite subset $S$ of the vertex set of $C$ such that $f(X)\subseteq C\cap\textnormal{Prob}(S)$. Define
@@
Since $S$ is finite and the stabilizer of each vertex in $S$ is finite, $F$ is a finite subset of $\Gamma$. Observe that $C$ is a $(\Gamma,\mathcal{F})$-complex $C$ of dimension at most $h(R)$, there exists an open cover $\{V_{0},\cdots,V_{h(R)}\}$ of $C$ as in [[GWY17], Lemma 4.1]. For each $i\in\{0,\cdots,h(R)\}$, and each $i$-simplex $\Delta$, we define $U_{i}:=f^{-1}(V_{i})$, $U_{i\Delta}:=f^{-1}(V_{i\Delta})$. Therefore, $\{U_{0},\cdots,U_{h(R)}\}$ is an open cover of $X$, and $U_{i}=\bigsqcup\limits_{\Delta\text{ an i-simplex }}U_{i\Delta}$. Indeed, $f(X)\subseteq C\cap\textnormal{Prob}(S)\subseteq\bigcup\limits_{i=0}^{h(R)}V_{i}$ implies that $X\subseteq f^{-1}(\bigcup\limits_{i=0}^{h(R)}V_{i})=\bigcup\limits_{i=0}^{h(R)}f^{-1}(V_{i})=\bigcup\limits_{i=0}^{h(R)}U_{i}$. By [[GWY17], Lemma 4.1], we know that $V_{i\Delta}$ and $V_{i\Delta^{\prime}}$ are distinct for different $i$-simplices, which implies that $f^{-1}(V_{i\Delta})$ and $f^{-1}(V_{i\Delta^{\prime}})$ are distinct, hence $U_{i\Delta}$ and $U_{i\Delta^{\prime}}$ are distinct.
Next, we claim that the following statement holds for any $i$ and $i$-simplex $\Delta$:
@@
As $x\in U_{i\Delta}$, $f(x)\in V_{i\Delta}$, which implies that $gf(x)\in V_{ig\Delta}$. Clearly, $gx\in f^{-1}(V_{i\Delta^{\prime}})$ implies $f(gx)\in V_{i\Delta^{\prime}}$. Observe that $d(V_{i\Delta^{\prime}},V_{ig\Delta})\geq\frac{1}{3}10^{-h(R)}$ when $\Delta^{\prime}\neq g\Delta$, but $d(f(gx),gf(x))<\frac{1}{3}10^{-h(R)}$, which forces $\Delta^{\prime}=g\Delta$. Hence, $gx\in U_{ig\Delta}$ as claimed.
Finally, we show that the action $\Gamma\curvearrowright X$ has dynamic asymptotic dimension function at most $h$. Suppose that $x\in U_{i}$ for some $i$ and $g=g_{n}\cdots g_{1}$ with $g_{k}\in E$ and $g_{k}\cdots g_{1}x\in U_{i}$, $\forall k\in\{1,\cdots,n\}$. If $x\in U_{i\Delta}$ for some $i$-simplex $\Delta$, then $gx\in U_{ig\Delta}$ by using the claim repeatedly, thus $f(gx)\in V_{ig\Delta}\cap f(X)\subseteq C\cap\textnormal{Prob}(S)$. Observe that $f(gx)$ and $f(x)$ are supported on vertices on $S$, which forces $gS\cap S\neq\varnothing$; thus $g$ belongs to the finite set $F$ By Definition 3.2, $\textnormal{dad}_{\Gamma\curvearrowright X}(R)\leq h(R)$, thus $\textnormal{dad}_{\Gamma\curvearrowright X}\leq h$. ∎
@@ 4. Dynamic asymptotic dimension growth for groupoids
In this section, we formulate the definition of dynamic asymptotic dimension growth for groupoids, which generalizes our notion for the case of group actions. Moreover, we will show that our notion is closely related to coarse geometry, the asymptotic dimension growth for countable discrete groups, and the amenability of groupoids.
@@ 4.1. The case of étale groupoids
We now focus on a class of étale groupoids, where it is easier to find conditions to derive amenability. Using the notion of length functions on étale groupoids, we may define the dynamic asymptotic dimension growth for étale groupoids. Furthermore, we will show that this notion is a generalization of the dynamic asymptotic dimension growth for group actions in this subsection.
@@
Furthermore, an étale groupoid enjoys the following properties.
@@
    continuous if it is continuous function with regard to the topology of $\mathcal{G}$.
@@
To obtain the amenability of a groupoid, we are mainly concerned about the coarse continuous length function. In the work of [MW20], Ma and Wu demonstrated the existence of coarse continuous length functions in the following case.
@@
Here is a typical example of the above theorem.
@@
$\qquad\qquad\qquad\qquad\qquad\qquad(\gamma x,\gamma,x)\mapsto\ell_{\Gamma}(\gamma)(1+\max\{g(\gamma x),g(x)\}).$
@@
We are mainly interested in étale groupoids with compact unit spaces, and we now define the dynamic asymptotic dimension growth for such groupoids.
@@
Let $\mathcal{G}$ be an étale groupoid with compact unit space, and let $\ell:\mathcal{G}\rightarrow[0,\infty)$ be a coarse continuous length function. For any $R>0$, let $\textnormal{dad}_{\mathcal{G}}(R)$ be the smallest number over all $m\in\mathbb{N}$ with the following property: for the open precompact subset $K={\{z\in\mathcal{G}\mid\ell(z)<R\}}$ of $\mathcal{G}$, there exists an open cover $\{U_{0},\cdots,U_{m}\}$ of $\mathcal{G}^{(0)}$ such that for each $i$, the set $\{g\in K\mid s(g),r(g)\in U_{i}\}$ is contained in a precompact subgroupoid of $\mathcal{G}$. We call $\textnormal{dad}_{\mathcal{G}}:\mathbb{R}^{+}\rightarrow\mathbb{N}$ the dynamic asymptotic dimension function of $\mathcal{G}$, and call the growth type of the function $\textnormal{dad}_{\mathcal{G}}$ the dynamic asymptotic dimension growth of $\mathcal{G}$.
@@
    If $\textnormal{dasdim}(\mathcal{G})=\infty$, then both sides of Equation 4.1 are infinite.
@@
    For each $i$, the subgroupoid $\mathcal{G}_{i}$ generated by the set
@@
The following lemma will be used in the proof of a key lemma in Section 5.
@@
To show that Definition 4.6 generalizes the dynamic asymptotic dimension growth for group actions in the sense of Definition 3.2, we consider the transformation groupoids induced by group actions on compact space $X$ in Example 4.5.
@@
Let $X$ be a compact Hausdorff space, and let $f:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a nondecreasing function. An action $\Gamma\curvearrowright X$ has dynamic asymptotic dimension function $f$ if and only if the corresponding transformation groupoid $\mathcal{G}=X\rtimes\Gamma$ has dynamic asymptotic dimesnion function $f$.
@@
We may assume that an action $\Gamma\curvearrowright X$ has dynamic asymptotic dimension function $f$, and its associated transformation groupoid has dynamic asymptotic dimension funnction $\textnormal{dad}_{\mathcal{G}}$. It suffices to show that $\textnormal{dad}_{\mathcal{G}}=f$. For any $R>0$, for the open precompact subset $K={\{(\gamma x,\gamma,x)\in\mathcal{G}\mid\ell(\gamma x,\gamma,x)<R\}}$ of $\mathcal{G}$. Also, $\ell_{\Gamma}(\gamma)=\ell(\gamma x,\gamma,x)<R$, thus $K$ has the form
@@
for a finite subset $E=\{\gamma\in\Gamma\mid\ell_{\Gamma}(\gamma)<R\}$ of $\Gamma$. Then there is an open cover $U_{0},\cdots,U_{f(R)}$ of $\mathcal{G}^{(0)}=X$ such that the set
@@
is finite for each $i\in\{0,\cdots,f(R)\}$. Therefore, the subgroupoid $\mathcal{G}_{i}$ generated by the set
@@
is contained in $\{(\gamma x,\gamma,x)\in\mathcal{G}\mid x\in X,\gamma\in F_{i}\}.$ Since $F_{i}$ is finite and $X$ is compact, the set $\{(\gamma x,\gamma,x)\in\mathcal{G}\mid x\in X,\gamma\in F_{i}\}$ is precompact, and thus $\mathcal{G}_{i}$ is precompact. Hence, $\textnormal{dad}_{\mathcal{G}}(R)\leq f(R)$.
On the other hand, for any $R>0$, for the finite subset $E=\{\gamma\in\Gamma\mid\ell_{\Gamma}(\gamma)<R\}$ of $\Gamma$, there is an open precompact subset $K=\{(\gamma x,\gamma,x)\in\mathcal{G}\mid x\in X,\gamma\in E\}$ of $\mathcal{G}$. Since $\mathcal{G}$ has the dynamic asymptotic dimension function $\textnormal{dad}_{\mathcal{G}}$, there is an open cover $\{U_{0},\cdots,U_{\textnormal{dad}_{\mathcal{G}}(R)}\}$ of $X$ such that for each $i\in\{0,\cdots,\textnormal{dad}_{\mathcal{G}}(R)\}$, the set
@@
is contained in $\mathcal{G}_{i}$, and is thus precompact, which implies that $F_{i}$ is finite. Clearly, $f(R)\leq\textnormal{dad}_{\mathcal{G}}(R)$, which forces $f(R)=\textnormal{dad}_{\mathcal{G}}(R)$. Hence, $\textnormal{dad}_{\mathcal{G}}=f$.
The converse can be completed by using the method analogous to that used above. ∎

### v2

@@ 3.2. Bartels-Lück-Reich condtions
Elements of $\textnormal{Prob}(V)$ can be written as $\mu=\sum\limits_{v\in V}t_{v}v$, where $\mu(v)=t_{v}$ in $[0,1]$ and $\sum\limits_{v\in V}t_{v}=1$. Write $\textnormal{Prob}_{n}(V)$ for the subset of $\textnormal{Prob}(V)$ consisting of measures supported on at most $n+1$ points. Say that a simplicial complex is a subspace $C$ of some $\bigcup\limits_{n=0}^{\infty}\textnormal{Prob}_{n}(V)$ such that $V\subseteq C$ and
$\left\{\sum\limits_{v\in V}t_{v}v\in\bigcup\limits_{n=0}^{\infty}\textnormal{Prob}_{n}(V)\bigg{|}t_{v}\neq 0\text{ for }v\in\textnormal{supp }(\mu)\right\}\in C.$
@@
Below we mainly focus on a simplicial complex equipped with a simplicial action, and let us first review this notion. Let $\Gamma$ be a discrete group, and let $C$ be a simplicial complex with the vertex set $V$. A simplicial action $g$ of $\Gamma$ on $C$ is an action that is induced from some action of $\Gamma$ on $V$ via the formula
A simplicial action is isometric.
The concept of an almost equivariant map from a topological space to a metric space is defined below.
@@
The next technical lemma will be used in the proof of Proposition 3.13 and Theorems 3.15 and 4.22.
@@
The underlying concept serves as the fundamental element of the Bartels-Lück-Reich conditions, which can be applied to families of finite subgroups or virtually cyclic subgroups.
@@
We now define the $h$-BLR action, which is close to the amenability of group actions.
@@
If $h\preceq x^{\alpha}(0<\alpha<\frac{1}{2})$, we can obtain an equivalent definition of the $h$-BLR action, which is similar to the definition of asymptotic dimension growth for a group.
@@
    For any $R>0$, there exists a nondecreasing function $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ such that for the finite subset $E=\{\gamma\in\Gamma\mid\ell_{\Gamma}(\gamma)<R\}$ of $\Gamma$, there exists a $\Gamma$-equivaraint open cover $\mathcal{U}$ of $X\times\Gamma$ such that the following hold.
@@
This proposition can be proved in a similar way as shown in [[GWY17], Proposition 4.5]. In the proof of $(1)$ implying $(2)$, there are no extra restrictions on the growth rate of $h$, while in the case of $(2)$ implying $(1)$, the function $h$ is required to grow at a rate slower than $\sqrt{x}$.
The following theorem shows that an $h$-BLR action has dynamic asymptotic dimension function at most $h$, which plays an important role in the proof Theorem 4.21.
@@
Let $\Gamma\curvearrowright X$ be an action with $X$ compact, and let $\mathcal{F}$ be the family of finite subgroups of $\Gamma$. Let $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a nondecreasing function. If $\Gamma\curvearrowright X$ is an $h$-BLR action for the family $\mathcal{F}$, then it has dynamic asymptotic dimension function at most $h$.
@@
We divide the proof into three steps. First, we construct a desired open cover of $X$. Assume that $\Gamma\curvearrowright X$ is an $h$-$BLR$ action for $\mathcal{F}$. Given $R>0$, there exists a nondecreasing function $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ such that for the finite subset $E=\{\gamma\in\Gamma\mid\ell(\gamma)<R\}$ of $\Gamma$ and $\varepsilon=\frac{1}{3}10^{-h(R)}$, there exist a $(\Gamma,\mathcal{F})$-complex $C$ of dimension at most $h(R)$ and a continuous $(E,\varepsilon)$-equivariant map $f:X\rightarrow C$. By Lemma 3.10, we may assume that there exists a finite subset $S$ of the vertex set of $C$ such that $f(X)\subseteq C\cap\textnormal{Prob}(S)$. Define
@@
Since $S$ is finite and the stabilizer of each vertex in $S$ is finite, $F$ is a finite subset of $\Gamma$. Observe that $C$ is a $(\Gamma,\mathcal{F})$-complex $C$ of dimension at most $h(R)$, there exists an open cover $\{V_{0},\cdots,V_{h(R)}\}$ of $C$ as in [[GWY17], Lemma 4.1]. For each $i\in\{0,\cdots,h(R)\}$, and each $i$-simplex $\Delta$, we define $U_{i}:=f^{-1}(V_{i})$, $U_{i\Delta}:=f^{-1}(V_{i\Delta})$. Therefore, $\{U_{0},\cdots,U_{h(R)}\}$ is an open cover of $X$, and $U_{i}=\bigsqcup\limits_{\Delta\text{ an i-simplex }}U_{i\Delta}$. Indeed, $f(X)\subseteq C\cap\textnormal{Prob}(S)\subseteq\bigcup\limits_{i=0}^{h(R)}V_{i}$, which implies that $X\subseteq f^{-1}(\bigcup\limits_{i=0}^{h(R)}V_{i})=\bigcup\limits_{i=0}^{h(R)}f^{-1}(V_{i})=\bigcup\limits_{i=0}^{h(R)}U_{i}$. By [[GWY17], Lemma 4.1], we know that $V_{i\Delta}$ and $V_{i\Delta^{\prime}}$ are distinct for different $i$-simplices, which implies that $f^{-1}(V_{i\Delta})$ and $f^{-1}(V_{i\Delta^{\prime}})$ are distinct, hence $U_{i\Delta}$ and $U_{i\Delta^{\prime}}$ are distinct.
Next, we claim that the following statement holds for each $i$ and $i$-simplex $\Delta$:
@@
Since $x\in U_{i\Delta}$, we have $f(x)\in V_{i\Delta}$, which implies that $gf(x)\in V_{ig\Delta}$. Clearly, $gx\in f^{-1}(V_{i\Delta^{\prime}})$ implies $f(gx)\in V_{i\Delta^{\prime}}$. Observe that $d(V_{i\Delta^{\prime}},V_{ig\Delta})\geq\frac{1}{3}10^{-h(R)}$ when $\Delta^{\prime}\neq g\Delta$, but $d(f(gx),gf(x))<\frac{1}{3}10^{-h(R)}$, which forces $\Delta^{\prime}=g\Delta$. Hence, $gx\in U_{ig\Delta}$ as claimed.
Finally, we show that the action $\Gamma\curvearrowright X$ has dynamic asymptotic dimension function at most $h$. Suppose that $x\in U_{i}$ for some $i$ and $g=g_{n}\cdots g_{1}$ with $g_{k}\in E$ and $g_{k}\cdots g_{1}x\in U_{i}$, $\forall k\in\{1,\cdots,n\}$. If $x\in U_{i\Delta}$ for some $i$-simplex $\Delta$, then $gx\in U_{ig\Delta}$ by using the claim repeatedly, thus $f(gx)\in V_{ig\Delta}\cap f(X)\subseteq C\cap\textnormal{Prob}(S)$. Observe that both $f(gx)$ and $f(x)$ are supported on vertices on $S$, which forces $gS\cap S\neq\varnothing$. Therefore, $g$ belongs to the finite set $F$. According to Definition 3.2, it follows that $\textnormal{dad}_{\Gamma\curvearrowright X}(R)\leq h(R)$. Hence, we conclude that $\textnormal{dad}_{\Gamma\curvearrowright X}\leq h$. ∎
@@ 4. Dynamic asymptotic dimension growth for groupoids
In this section, we formulate the definition of dynamic asymptotic dimension growth for étale groupoids, which generalizes our notion for actions of discrete groups on compact spaces. Moreover, we will examine the relation between our notion with the asymptotic dimension growth. In coarse geometry, we show that the asymptotic dimension growth for a discrete metric space of bounded geometry $X$ is equivalent to the dynamic asymptotic dimension growth for its corresponding coarse groupoid $G(X)$, thus implying that the groupoid $G(X)$ with subexponential dynamic asymptotic dimension growth is amenable. Using this result, we establish the equivalence between the asymptotic dimension growth for a countable discrete group and the dynamic asymptotic dimension growth for its action on some compact Hausdorff space. Consequently, we obtain that a countable discrete group with subexponential asymptotic dimension growth admits a free, amenable, minimal action on the Cantor set which has at most subexponential dynamic asymptotic dimension growth. This allows us to provide many concrete examples for our notion.
@@ 4.1. The case of étale groupoids
We now focus on étale groupoids, from which it is easier to find sufficient conditions for the amenability of groupoids. In this subsection, we introduce the dynamic asymptotic dimension growth for étale groupoids. We first define the dynamic asymptotic dimension growth for étale groupoids with compact unit spaces by using length functions on these groupoids. Further, borrowing the notion of Alexandrov one-point compactification for groupoids defined in [KKLRU21], we can define the dynamic asymptotic dimension growth for étale groupoids with non-compact unit spaces. Finally, we show that this notion is a generalization of the dynamic asymptotic dimension growth for actions of discrete groups on compact spaces.
@@
The length functions on étale groupoids enjoy the following properties.
@@
    continuous if it is a continuous function with regard to the topology of $\mathcal{G}$.
@@
To obtain the amenability of groupoids, we are mainly concerned about coarse continuous length functions. The existence of coarse continuous length functions was examined by Ma and Wu in [MW20] in the following case.
@@
The following transformation groupoid is a typical example of the above theorem, which plays an important role in the proof of Lemma 4.12.
@@
$\quad\qquad\qquad\qquad\qquad\qquad(\gamma x,\gamma,x)\mapsto\ell_{\Gamma}(\gamma)(1+\max\{g(\gamma x),g(x)\}).$
@@
By using coarse continuous length functions on étale groupoids with compact unit spaces, we can define the dynamic asymptotic dimension growth for these groupoids.
@@
Let $\mathcal{G}$ be an étale groupoid with compact unit space $\mathcal{G}^{(0)}$, and let $\ell:\mathcal{G}\rightarrow[0,\infty)$ be a coarse continuous length function. For any $R>0$, let $\textnormal{dad}_{\mathcal{G}}(R)$ be the smallest number over all $m\in\mathbb{N}$ with the following property: for the open precompact subset $K={\{z\in\mathcal{G}\mid\ell(z)<R\}}$ of $\mathcal{G}$, there exists an open cover $\{U_{0},\cdots,U_{m}\}$ of $\mathcal{G}^{(0)}$ such that for each $i$, the set $\{g\in K\mid s(g),r(g)\in U_{i}\}$ is contained in a precompact subgroupoid of $\mathcal{G}$. We call $\textnormal{dad}_{\mathcal{G}}:\mathbb{R}^{+}\rightarrow\mathbb{N}$ the dynamic asymptotic dimension function of $\mathcal{G}$, and call the growth type of the function $\textnormal{dad}_{\mathcal{G}}$ the dynamic asymptotic dimension growth of $\mathcal{G}$. If no such function $\textnormal{dad}_{\mathcal{G}}$ exists, we say that $\mathcal{G}$ has infinite dynamic asymptotic dimension.
If $\mathcal{G}$ is an étale groupoid with non-compact unit space, the following definition provides a suitable way to reduce the unit space to the compact case.
Definition 4.7 ([KKLRU21]).
Let $\mathcal{G}$ be an étale groupoid with locally compact Hausdorff space of units. Then the Alexandrov groupoid $\mathcal{G}^{+}$ is the set $\mathcal{G}\cup(\mathcal{G}^{(0)})^{+}$ with the topology determined by specifying that the inclusions $\mathcal{G}$, $(\mathcal{G}^{(0)})^{+}\subseteq\mathcal{G}^{+}$ are open, and the groupoid structure extending the groupoid structure of $\mathcal{G}$ and making $\infty$ a unit.
Observe that $\mathcal{G}^{+}$ is an étale groupoid whose unit space is the compact Hausdorff space $(\mathcal{G}^{(0)})^{+}$ by taking Alexandrov one-point compactification of $\mathcal{G}^{(0)}$. The next proposition demonstrated that the original groupoid $\mathcal{G}$ shares the same dynamic asymptotic dimension with its Alexandrov groupoid $\mathcal{G}^{+}$.
Proposition 4.8 ([CDGHV24]).
Let $\mathcal{G}$ be a locally compact, Hausdorff and étale groupoid. Then $\textnormal{dad}(\mathcal{G})=\textnormal{dad}(\mathcal{G}^{+})$.
For groupoids with non-compact unit spaces, we apply the above proposition to Definition 4.6, and let $\textnormal{dad}_{\mathcal{G}}(R):=\textnormal{dad}_{\mathcal{G}^{+}}(R)$, for any $R>0$. In this way, we can define the dynamic asymptotic dimension growth for groupoids with non-compact unit spaces.
Definition 4.9.
Let $\mathcal{G}$ be an étale groupoid with non-compact unit space $\mathcal{G}^{(0)}$, and let $\mathcal{G}^{+}$ be the Alexandrov groupoid of $\mathcal{G}$. We define the dynamic asymptotic dimension function of $\mathcal{G}$ as $\textnormal{dad}_{\mathcal{G}}:=\textnormal{dad}_{\mathcal{G}^{+}}$, and call the growth type of the function $\textnormal{dad}_{\mathcal{G}}$ the dynamic asymptotic dimension growth of $\mathcal{G}$.
@@
    If $\textnormal{dasdim}(\mathcal{G})=\infty$, then both sides of equation (4.1) are infinite.
@@
    For each $i$, the subgroupoid $\mathcal{G}_{i}=\langle K\cap\mathcal{G}|_{U_{i}}\rangle$ generated by the set
@@
The following lemma will be used in the proof of Lemma 5.6.
@@
Using the length function on the transformation groupoid constructed in Example 4.5, we prove that the group action $\Gamma\curvearrowright X$ shares the same dynamic asymptotic dimension function with its corresponding transformation groupoid $X\rtimes\Gamma$. This means that Definition 4.6 generalizes the dynamic asymptotic dimension growth for group actions in the sense of Definition 3.2.
@@
Let $X$ be a compact Hausdorff space, and let $f:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a nondecreasing function. An action $\Gamma\curvearrowright X$ has dynamic asymptotic dimension function $f$ if and only if the corresponding transformation groupoid $\mathcal{G}=X\rtimes\Gamma$ has dynamic asymptotic dimension function $f$.
@@
We may assume that an action $\Gamma\curvearrowright X$ has dynamic asymptotic dimension function $\textnormal{dad}_{\Gamma\curvearrowright X}$, and its associated transformation groupoid has dynamic asymptotic dimension function $\textnormal{dad}_{\mathcal{G}}$. It thus suffices to show that $\textnormal{dad}_{\mathcal{G}}=\textnormal{dad}_{\Gamma\curvearrowright X}$. First, we prove that $\textnormal{dad}_{\mathcal{G}}\leq\textnormal{dad}_{\Gamma\curvearrowright X}$. For any $R>0$, for the open precompact subset $K={\{(\gamma x,\gamma,x)\in\mathcal{G}\mid\ell(\gamma x,\gamma,x)<R\}}$ of $\mathcal{G}$. Then $\ell_{\Gamma}(\gamma)=\ell(\gamma x,\gamma,x)<R$, thus $K$ has the form
@@
for a finite subset $E=\{\gamma\in\Gamma\mid\ell_{\Gamma}(\gamma)<R\}$ of $\Gamma$. Since the action has dynamic asymptotic dimension function $\textnormal{dad}_{\Gamma\curvearrowright X}$, there is an open cover $U_{0},\cdots,U_{\textnormal{dad}_{\Gamma\curvearrowright X}(R)}$ of $\mathcal{G}^{(0)}=X$ such that the set
@@
is finite for each $i\in\{0,\cdots,\textnormal{dad}_{\Gamma\curvearrowright X}(R)\}$. Therefore, the subgroupoid $\mathcal{G}_{i}$ generated by the set
@@
is contained in $\{(\gamma x,\gamma,x)\in\mathcal{G}\mid x\in X,\gamma\in F_{i}\}.$ Since $F_{i}$ is finite and $X$ is compact, the set $\{(\gamma x,\gamma,x)\in\mathcal{G}\mid x\in X,\gamma\in F_{i}\}$ is precompact, and thus $\mathcal{G}_{i}$ is precompact. Hence, $\textnormal{dad}_{\mathcal{G}}(R)\leq\textnormal{dad}_{\Gamma\curvearrowright X}(R)$, which implies that $\textnormal{dad}_{\mathcal{G}}\leq\textnormal{dad}_{\Gamma\curvearrowright X}$.
On the other hand, we show that $\textnormal{dad}_{\Gamma\curvearrowright X}\leq\textnormal{dad}_{\mathcal{G}}$. For any $R>0$, for the finite subset $E=\{\gamma\in\Gamma\mid\ell_{\Gamma}(\gamma)<R\}$ of $\Gamma$, there is an open precompact subset $K=\{(\gamma x,\gamma,x)\in\mathcal{G}\mid x\in X,\gamma\in E\}$ of $\mathcal{G}$. Since $\mathcal{G}$ has the dynamic asymptotic dimension function $\textnormal{dad}_{\mathcal{G}}$, there is an open cover $\{U_{0},\cdots,U_{\textnormal{dad}_{\mathcal{G}}(R)}\}$ of $X$ such that for each $i\in\{0,\cdots,\textnormal{dad}_{\mathcal{G}}(R)\}$, the set
@@
is contained in $\mathcal{G}_{i}$, and is thus precompact, which implies that $F_{i}$ is finite. Clearly, $\textnormal{dad}_{\Gamma\curvearrowright X}(R)\leq\textnormal{dad}_{\mathcal{G}}(R)$, which forces $\textnormal{dad}_{\Gamma\curvearrowright X}\leq\textnormal{dad}_{\mathcal{G}}$. Hence, $\textnormal{dad}_{\Gamma\curvearrowright X}=\textnormal{dad}_{\mathcal{G}}$, namely, $\textnormal{dad}_{\Gamma\curvearrowright X}=f$ if and only if $\textnormal{dad}_{\mathcal{G}}=f$. ∎
Remark 4.13.
Using the above notation, let $\mathcal{G}^{+}$ be the Alexandrov groupoid of $\mathcal{G}$, and let $X_{+}$ be the one-point compactification of $X$. It is easy to see that the action of $\Gamma$ on $X_{+}$ given by
$\gamma\cdot x=\begin{cases}\gamma\cdot x&\text{ if }x\in X\\ \infty&\text{ if }x=\infty\end{cases}$
is continuous. Then we have
$\mathcal{G}^{+}=\{(\gamma,x):\gamma\in\Gamma,x\in X\}\cup\{(e,\infty)\}\text{ and }X_{+}\rtimes\Gamma=\{(\gamma,x):\gamma\in\Gamma,x\in X_{+}\}.$
It is easy to verify that the Alexandrov groupoid $\mathcal{G}^{+}$ is not identical to the transformation groupoid $X_{+}\rtimes\Gamma$; it is merely its wide and open subgroupoid [[CDGHV24], Example 3.5]. Therefore, using the Alexandrov construction, the above lemma does not extend to the non-compact Hausdorff space $X$.